OpenStudy (anonymous):

10 in. = ________ft. {give the answer in a fraction}

jigglypuff314 (jigglypuff314):

there is 12 inches in 1 foot so to convert inches to feet (number of inches) divided by 12 = (number of feet)

jigglypuff314 (jigglypuff314):

replace (number of inches) with 10 so 10 divided by 12 is the fraction :)

OpenStudy (doc.brown):

Or do it properly.\[10in\times\frac{ft}{12in}\]\[10\cancel{in}\times\frac{ft}{12\cancel{in}}\]\[=\frac{10}{12}ft\]You can probably reduce that.

OpenStudy (anonymous):

Thank you so much @doc.brown . Can u please tell me what is the the median of the following numbers 250 270 280 280 320?